OpenStudy (anonymous):

the value of "n" in n/40 lie between 1/5 and 1/4 , determine the possible value .. how would I get this answer? & what is the answer?!!!!!!!!! :)

OpenStudy (anonymous):

You'd need to convert 1/5 and 1/4 to 40ths: 1/5=8/40, 1/4=10/40. If n/40 lies between 1/5 and 1/4, then n would have to be 9 to make it 9/40.